# Topology of Discrete Quantum Feedback Control

### v2

@@ VIII.1 Boundary-localized steady state
Complete positivity and trace preservation ensure that a quantum channel gives a map between density matrices, which leads to the existence of a steady state with unit eigenvalue [102]. Consequently, the topological Maxwell’s demons discussed in Sec. IV and VII (except for those in Sec. VII.4 and VII.5) exhibit boundary-localized steady states under the OBC. This is in stark contrast to topological phases described by non-Hermitian Hamiltonians, where the conservation of the norm of a state is not guaranteed and therefore a steady state does not exist in general. Although those boundary-localized steady states in the topological feedback control appear to be closely related to the non-Hermitian skin effect, we can construct a model that shows the non-Hermitian skin effect with no localized steady state (see Appendix F). In this sense, the emergence of boundary-localized steady states is a special feature of quantum channels and should be distinguished from the non-Hermitian skin effect. We note that a similar phenomenon is observed in classical stochastic processes where the probability conservation is satisfied unlike general non-Hermitian systems [103].
@@ VIII.2 Breakdown of conventional bulk-boundary correspondence in non-Hermitian skin effects
The bulk-boundary correspondence in non-Hermitian systems with nontrivial point-gap topology states that a nonzero winding number of the eigenspectrum under the PBC leads to boundary-localized skin modes under the OBC whose eigenvalues generically reside inside the region delimited by the PBC spectrum [72]. However, eigenvalues of the boundary-localized steady states discussed in Sec. VIII.1 must be unity by definition and do not belong to the internal region enclosed by the PBC spectrum (see e.g., Figs. 6, 8, and 13). More strikingly, it is possible that some eigenvalues under the OBC lie outside the region enclosed by the PBC spectrum (see the case of the QND demon shown in Fig. 16). These facts indicate that the conventional bulk-boundary correspondence in non-Hermitian skin effects does not hold in topological feedback control.
@@ VIII.3 Topological distinction between steady states and transient dynamics
The complete positivity also impacts the topological classification. As discussed in Sec. V.2, the complete positivity (or Hermiticity preservation following therefrom) imposes the modular conjugation symmetry on a quantum channel. While some symmetry sectors of a quantum channel do not necessarily satisfy the modular conjugation symmetry (as shown in Sec. VI.3), a sector that includes a steady state must satisfy it since the steady-state density matrix must be Hermitian [93]. Thus, symmetry classes allowed for a steady-state sector of a quantum channel are restricted to those three classes (AI, AI $+$ psH+, and AI $+$ psH-) that include the modular conjugation symmetry. This means that topological classification for boundary-localized steady states in Sec. VIII.1 is more restrictive than the general topological classification of non-Hermitian systems due to the complete positivity.
If transient dynamics are considered, topological phenomena that are prohibited in steady states due to complete positivity become possible. An example is given by the $\mathbb{Z}_{2}$ demon in Sec. VII.4. As shown in Fig. 18, a spin-diagonal sector $\mathcal{E}_{\uparrow\uparrow}$ that includes a steady state is topologically trivial, whereas an off-diagonal sector $\mathcal{E}_{\uparrow\downarrow}$ displays nontrivial point-gap topology owing to the absence of the modular conjugation symmetry. The nontrivial topology in the off-diagonal sector manifests itself in the transient dynamics of the transverse components of the spin density, which vanish in the steady state (see Fig. 19). The topological distinction between the steady state and transient dynamics thus enriches topological phenomena in quantum feedback control.
@@ IX.1 Formalism
Then, given an initial density matrix $\hat{\rho}_{\mathrm{i}}$, the density matrix after feedback control is given by
@@
where $p_{m}=\mathrm{Tr}[\hat{M}_{m}^{\dagger}\hat{M}_{m}\hat{\rho}_{\mathrm{i}}]$ is the probability of outcome $m$ for the measurement operator $\hat{M}_{m}$. Thus, the feedback control is described by a quantum channel
@@
IX.2 Dissipative Maxwell’s demon
As an example of dissipative feedback control described by a CPTP map (173), we consider a chiral Maxwell’s demon (see Sec. IV) subject to dissipation. We consider a spinless particle on a 1D lattice with $L$ sites, and perform projective position measurement for this particle as in Sec. IV. We then perform a feedback operation governed by the Lindblad equation
@@ IX.2 Dissipative Maxwell’s demon
where $\mathcal{L}_{m}$ is the generator (Lindbladian) conditioned on the measurement outcome, $H_{m}$ is a feedback Hamiltonian given in Sec. IV.1, and $\hat{L}_{m,j}$ is the Lindblad operator acting on a local region around site $j$. Then, a CPTP map for the feedback operation is given by
@@
where $\tau$ is the duration of feedback. Using the Lindblad equation (180), we can consider various dissipative effects on feedback control such as noise during feedback and a coupling to a reservoir. In particular, if the system is in contact with a heat bath, the topological Maxwell’s demon can rectify thermal fluctuations as well as quantum fluctuations to achieve unidirectional transport of a particle. However, we note that the Hamiltonian and the Lindblad operators in Eq. (180) should be chosen to be local to ensure the locality of the Kraus operators and to allow truncation of the Bloch matrix into a finite-dimensional matrix (see Sec. III.2). The locality-preserving Lindblad equation can be derived if the correlation time of the environment is sufficiently short [104].
@@
the Bloch matrix of the quantun channel is given by
@@
Here we note that
@@
where $p(j|m)$ is the probability of a particle being found at site $j$ after the feedback control with outcome $m$. Then, the eigenvalue $\xi_{0}(k)$ is written as
@@
which takes the same form as in the case with unitary feedback control [see Eq. (43)]. Thus, regardless of whether the feedback control is unitary or not, the eigenvalues of the quantum channel and its topology are completely determined from the probability distribution $p(j|m)$ of the position after the feedback.
To demonstrate the robustness of the chiral Maxwell’s demon against dissipation during feedback control, we consider the Lindblad operators
$L_{m,j}^{(\mathrm{d})}=$ $\sqrt{\gamma_{\mathrm{d}}}\ket{j}\bra{j},$ (188)
$L_{m,j}^{(\mathrm{L})}=$ $\sqrt{\gamma_{\mathrm{L}}}\ket{j-1}\bra{j},$ (189)
$L_{m,j}^{(\mathrm{R})}=$ $\sqrt{\gamma_{\mathrm{R}}}\ket{j+1}\bra{j},$ (190)
which respectively describe dephasing, stochastic hopping to the left direction, and stochastic hopping to the right direction [see Fig. 21(a) for a schematic illustration]. Figure 21(b) shows the eigenspectrum of the quantum channel for the chiral Maxwell’s demon subject to dissipation. As can be seen from the figure, the loop structure of the eigenspectrum under the PBC is robust against dissipation, leading to the non-Hermitian skin effect under the OBC. This result indicates that the chiral Maxwell’s demon can rectify the noise induced by dissipation as well as quantum fluctuations induced by unitary evolution, and achieve unidirectional transport that is robust against noise and decoherence.
Figure 21: (a) Schematic illustration of a chiral Maxwell’s demon subject to dissipation. (b) Eigenspectrum of a quantum channel for a chiral Maxwell’s demon subject to dissipation. The parameters are set to be $J=1,V=10,\gamma_{\mathrm{d}}=0.5,\gamma_{\mathrm{L}}=\gamma_{\mathrm{R}}=0.25$, and $\tau=1$. A blue curve and blue dots at the origin constitute the eigenspectrum under the PBC in the limit of infinite system size, and orange dots show the eigenspectrum under the OBC with the system size $L=30$.
Figure 22: (a) Schematic illustration of topological feedback control with engineered dissipation. (b) Eigenspectrum of a quantum channel for topological feedback control with engineered dissipation. The parameters are set to be $J=1,\gamma_{\mathrm{d}}=5$, and $\tau=1$. A blue curve and blue dots at the origin constitue the eigenspectrum under the PBC in the limit of infinite system size, and orange dots show the eigenspectrum under the OBC with the system size $L=30$.
@@ IX.3 Feedback control with engineered dissipation
Dissipative feedback control not only shows the stability of topological feedback control against noise, but also can be exploited to utilize dissipation as an essential tool for feedback control. Such a possibility can be envisaged since recent developments in quantum technologies with atoms, ions, and photons have enabled one to realize engineered dissipation [105]. For example, localized dissipation can be introduced in ultracold atoms, where the strength of dissipation is controlled by laser fields [106, 107, 108, 109, 110]. Motivated by the possibility of dissipation engineering, we propose a topological feedback control that is enabled by dissipation.
The setup of feedback control with engineered dissipation is as follows. We consider a spinless particle on a 1D lattice, perform projective position measurement on it, and obtain a measurement outcome $m$. Then, we let the particle undergo dissipative time evolution under the Lindblad equation (180). Here, we assume that the Hamiltonian does not depend on the measurement outcome $m$ and is given by a simple tight-binding model
$H_{m}=H:=-J\sum_{j}(\ket{j}\bra{j+1}+\mathrm{H.c.})$ (191)
with $J>0$ being the hopping amplitude. The dependence of the conditioned Lindblad equation (180) on the measurement outcome is encoded in the Lindblad operator, where engineered disipation is described by
$L_{m,j-1}=\delta_{m,j}\sqrt{\gamma_{\mathrm{d}}}\ket{j-1}\bra{j-1}.$ | . (192)
In other words, we introduce localized dissipation to the left site of the particle [see Fig. 22(a) for a schematic illustration]. Here we exploit the continuous quantum Zeno effect [111] in the strong-dissipation regime $\gamma_{\mathrm{d}}\gg J$ to suppress the hopping of the particle to the site that is subject to strong dissipation. Because of this feedback control of dissipation, hopping of a particle to the right is facilitated and chiral transport is eventually achieved.
The eigenspectrum of a quantum channel for this dissipative feedback control is shown in Fig. 22(b). The eigenspectrum under the PBC forms a loop and that under the OBC displays the non-Hermitian skin effect. Here, the loop structure of the eigenspectrum under the PBC is realized due to the dissipation, since the Hamiltonian (191) does not contain a wall potential. Thus, the topological feedback control, and the associated chiral transport, are achieved by dissipation.
@@ X Conclusion and outlook
In this paper, we have presented a class of dynamical topological phases realized with feedback control, which are chracterized by topology of quantum channels. The nontrivial point-gap topology of a quantum channel chracterized by a winding number ensures robust chiral tranport, which is analogous to the chiral edge current in the quantum Hall effect, and leads to the emergence of the non-Hermitian skin effect under the open boundary condition. We have also provided a symmetry classification of quantum feedback control, identified ten classes compatible with projective measurement, and constructed examples of feedback control in each symmetry class. This general framework of symmetry and topology of quantum feedback control provides a guiding princple for designing stable feedback control of quantum systems since topology is insensitive to small deformation. Also, feedback control offers a unique way of control of dynamical topological phases of matter distinct from other well-known classes such as periodically driven systems [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], non-Hermitian systems [32, 33, 34, 35, 36, 37, 38], and open quantum systems subject to engineered dissipation [64, 65].
@@
Topological feedback control is expected to be realized in various experimental platforms. For example, the past decade has witnessed remarkable progress in site-resolved measurement techniques and control of ultracold atoms in optical lattices [112, 113, 114]. More recently, atoms in optical tweezer arrays have emerged as a highly-controllable quantum simulator in which individual atoms can be measured and addressed in a programmable manner [115, 116, 117, 118, 119, 50, 120]. The high-precision position measurement and subsequent local control of a Hamiltonian for topological feedback control can be achieved by these state-of-the-art techniques. In addition to these promising platforms, quantum feedback control has widely been used in photonic systems [42, 58], levitated nanoparticles [44, 45, 46], and superconducting qubits [43, 60, 61, 62]. Given the rapid development in quantum control, topological feedback control in various experimental platforms will expand the scope of topological phases of matter.
@@
We conclude this paper by discussing future perspectives and open problems. First, while we have mostly focused on feedback control with projective measurement, the present framework for topology of quantum channels is also applicable to general quantum measurement, and therefore it is natural to ask whether non-projective measurement can be harnessed to construct topological feedback control. In particular, since the symmetry classification in Sec. V.3 is based on the spectral property of projective measurement channels, BL symmetry classes other than those in Table 2 are not prohibited for general quantum measurement. Since a quantum channel in a symmetry class beyond the ten-fold way necessarily has a pair of eigenvalues ±$\pm 1$ (if one assumes the existence of a steady state), its eigenvalue spectrum is reminiscent of that of dissipative time crystals [33, 121], where the eigenvalue $-1$ implies oscillating components under repeated application of a quantum channel.
@@
Second, as discussed in Sec. VIII, a crucial distinction between quantum channels and general non-Hermitian matrices is complete positivity and trace preservation (CPTP), which should be satisfied by any time evolution of the density matrix of a quantum system [47]. Although the topological classification of non-Hermitian matrices without the CPTP condition has been obtained [34, 35], it is unclear whether the CPTP condition alters the topological classification. For example, in Sec. VI.2, we have shown that the CPTP condition places a nontrivial constraint on the Bloch matrix with transposition symmetry, which turns out to be characterized by an integer topological invariant rather than a $\mathbb{Z}_{2}$ invariant. It merits further study to investigate the generality of this phenomenon. From a broader perspective, this problem may be related to a classification of time evolution of open quantum systems described by CPTP maps.
@@
Last but not least, topological feedback control suggests a novel connection between topology and Maxwell’s demon. For example, an experimental realization of Maxwell’s demon in Ref. [54] demonstrated chiral transport of a particle due to feedback control, in close analogy with the chiral Maxwell’s demon in Sec. IV (see also Ref. [82]). Maxwell’s demon has been of long-standing interest in thermodynamics and statistical physics, and an information-theoretic understanding of the role of Maxwell’s demon has led to the integration of thermodynamics and information theory [53]. Investigating topology of feedback control may thus shed light on a hitherto unnoticed role of topology in thermodynamics.

### v3

@@ VIII.1 Boundary-localized steady state
Complete positivity and trace preservation ensure that a quantum channel gives a map between density matrices, which leads to the existence of a steady state with the unit eigenvalue [102]. Consequently, the topological Maxwell demons discussed in Secs. IV and VII (except for those in Secs. VII.4 and VII.5) exhibit boundary-localized steady states under the OBC. This feature is in stark contrast to topological phases described by non-Hermitian Hamiltonians, where the conservation of the norm of a state is not guaranteed and therefore a steady state does not exist, in general. Although those boundary-localized steady states in topological feedback control appear to be closely related to the non-Hermitian skin effect, we can construct a model that shows the non-Hermitian skin effect with no localized steady state (see Appendix F). In this sense, the emergence of boundary-localized steady states is a special feature of quantum channels and should be distinguished from the non-Hermitian skin effect. We note that a similar phenomenon is observed in classical stochastic processes where the probability conservation is satisfied, unlike general non-Hermitian systems [103].
@@ VIII.2 Breakdown of conventional bulk-boundary correspondence in non-Hermitian skin effects
The bulk-boundary correspondence in non-Hermitian systems with nontrivial point-gap topology states that a nonzero winding number of the eigenspectrum under the PBC leads to boundary-localized skin modes under the OBC whose eigenvalues generically reside inside the region delimited by the PBC spectrum [72]. However, eigenvalues of the boundary-localized steady states discussed in Sec. VIII.1 must be unity by definition, and they do not belong to the internal region enclosed by the PBC spectrum (see, e.g., Figs. 6, 8, and 13). More strikingly, it is possible that some eigenvalues under the OBC lie outside the region enclosed by the PBC spectrum (see the case of the QND demon shown in Fig. 16). These facts indicate that the conventional bulk-boundary correspondence in non-Hermitian skin effects does not hold in topological feedback control.
@@ VIII.3 Topological distinction between steady states and transient dynamics
The complete positivity also impacts the topological classification. As discussed in Sec. V.2, the complete positivity (or the Hermiticity preservation that follows) imposes the modular conjugation symmetry on a quantum channel. While some symmetry sectors of a quantum channel do not necessarily satisfy the modular conjugation symmetry (as shown in Sec. VI.3), a sector that includes a steady state must satisfy it since the steady-state density matrix must be Hermitian [93]. Thus, allowed symmetry classes for a steady-state sector of a quantum channel are restricted to those three classes (AI, AI $+$ psH+, and AI $+$ psH-), which include modular conjugation symmetry. Thus, the topological classification for boundary-localized steady states in Sec. VIII.1 is more restrictive than the general topological classification of non-Hermitian systems due to the complete positivity.
If transient dynamics are considered, topological phenomena that are prohibited in steady states due to complete positivity become possible. An example is given by the $\mathbb{Z}_{2}$ demon in Sec. VII.4. As shown in Fig. 18, a spin-diagonal sector $\mathcal{E}_{\uparrow\uparrow}$ that includes a steady state is topologically trivial, whereas an off-diagonal sector $\mathcal{E}_{\uparrow\downarrow}$ displays nontrivial point-gap topology owing to the absence of modular conjugation symmetry. The nontrivial topology in the off-diagonal sector manifests itself in the transient dynamics of the transverse components of the spin density, which vanish in the steady state (see Fig. 19). The topological distinction between the steady state and transient dynamics thus enriches topological phenomena in quantum feedback control.
@@ IX.1 Formalism
Then, given an initial density matrix $\hat{\rho}_{\mathrm{i}}$, the density matrix after feedback control is
@@
where $p_{m}=\mathrm{Tr}[\hat{M}_{m}^{\dagger}\hat{M}_{m}\hat{\rho}_{\mathrm{i}}]$ is the probability of outcome $m$ for the measurement operator $\hat{M}_{m}$. Thus, feedback control is described by a quantum channel
@@
IX.2 Dissipative Maxwell demon
As an example of dissipative feedback control described by a CPTP map, Eq. (173), we consider a chiral Maxwell demon (see Sec. IV) subject to dissipation. We consider a spinless particle on a 1D lattice with $L$ sites and perform a projective position measurement for this particle as in Sec. IV. We then perform a feedback operation governed by the Lindblad equation
@@ IX.2 Dissipative Maxwell demon
where $\mathcal{L}_{m}$ is the generator (Lindbladian) conditioned on the measurement outcome, $H_{m}$ is the feedback Hamiltonian given in Sec. IV.1, and $\hat{L}_{m,j}$ is the Lindblad operator acting on a local region around site $j$. Then, a CPTP map for the feedback operation is given by
@@
where $\tau$ is the duration of feedback. Using the Lindblad equation (180), we can consider various dissipative effects on feedback control such as noise during feedback and a coupling to a reservoir. In particular, if the system is in contact with a heat bath, the topological Maxwell demon can rectify thermal fluctuations as well as quantum fluctuations to achieve unidirectional transport of a particle. However, we note that the Hamiltonian and the Lindblad operators in Eq. (180) should be chosen to be local to ensure the locality of the Kraus operators and to allow truncation of the Bloch matrix into a finite-dimensional matrix (see Sec. III.2). The locality-preserving Lindblad equation can be derived if the correlation time of the environment is sufficiently short [104].
@@
the Bloch matrix of the quantum channel is given by
@@
Here, we note that
@@
where $p(j|m)$ is the probability of a particle being found at site $j$ after feedback control with outcome $m$. Then, the eigenvalue $\xi_{0}(k)$ is written as
@@
which takes the same form as in the case with unitary feedback control [see Eq. (43)]. Thus, regardless of whether feedback control is unitary or not, the eigenvalues of the quantum channel and its topology are completely determined by the probability distribution $p(j|m)$ of the position after the feedback.
To demonstrate the robustness of the chiral Maxwell demon against dissipation during feedback control, we consider the Lindblad operators
$\hat{L}_{m,j}^{(\mathrm{d})}=$ $\sqrt{\gamma_{\mathrm{d}}}\ket{j}\bra{j},$ (188)
$\hat{L}_{m,j}^{(\mathrm{L})}=$ $\sqrt{\gamma_{\mathrm{L}}}\ket{j-1}\bra{j},$ (189)
$\hat{L}_{m,j}^{(\mathrm{R})}=$ $\sqrt{\gamma_{\mathrm{R}}}\ket{j+1}\bra{j},$ (190)
which, respectively, describe dephasing, stochastic hopping to the left, and stochastic hopping to the right [see Fig. 21(a) for a schematic illustration]. Figure 21(b) shows the eigenspectrum of the quantum channel for the chiral Maxwell demon subject to dissipation. As can be seen from the figure, the loop structure of the eigenspectrum under the PBC is robust against dissipation, leading to the non-Hermitian skin effect under the OBC. This result indicates that the chiral Maxwell demon can rectify the noise induced by dissipation as well as quantum fluctuations induced by unitary evolution and achieve unidirectional transport that is robust against noise and decoherence.
Figure 21: (a) Schematic illustration of a chiral Maxwell demon subject to dissipation. (b) Eigenspectrum of a quantum channel for a chiral Maxwell demon subject to dissipation. The parameters are set to $J=1,V=10,\gamma_{\mathrm{d}}=0.5,\gamma_{\mathrm{L}}=\gamma_{\mathrm{R}}=0.25$, and $\tau=1$. A blue curve and blue dots at the origin constitute the eigenspectrum under the PBC in the limit of infinite system size, and orange dots show the eigenspectrum under the OBC with system size $L=30$.
Figure 22: (a) Schematic illustration of topological feedback control with engineered dissipation. (b) Eigenspectrum of a quantum channel for topological feedback control with engineered dissipation. The parameters are set to $J=1,\gamma_{\mathrm{d}}=5$, and $\tau=1$. A blue curve and blue dots at the origin constitute the eigenspectrum under the PBC in the limit of infinite system size, and orange dots show the eigenspectrum under the OBC with system size $L=30$.
@@ IX.3 Feedback control with engineered dissipation
Dissipative feedback control not only shows the stability of topological feedback control against noise, but it can also be exploited to utilize dissipation as an essential tool for feedback control. Such a possibility can be envisaged since recent developments in quantum technologies with atoms, ions, and photons have enabled one to realize engineered dissipation [105]. For example, localized dissipation can be introduced in ultracold atoms, where the strength of dissipation is controlled by laser fields [106, 107, 108, 109, 110]. Motivated by the possibility of dissipation engineering, we propose a topological feedback control that is enabled by dissipation.
The setup of feedback control with engineered dissipation is as follows. We consider a spinless particle on a 1D lattice, perform a projective position measurement on it, and obtain a measurement outcome $m$. Next, we let the particle undergo dissipative time evolution under the Lindblad equation (180). Here, we assume that the Hamiltonian does not depend on the measurement outcome $m$ and that it is given by a simple tight-binding model
$\hat{H}_{m}=\hat{H}:=-J\sum_{j}(\ket{j}\bra{j+1}+\mathrm{H.c.})$ (191)
with $J>0$ being the hopping amplitude. The dependence of the conditioned Lindblad equation (180) on the measurement outcome is encoded in the Lindblad operator, where engineered dissipation is described by
$\hat{L}_{m,j-1}=\delta_{m,j}\sqrt{\gamma_{\mathrm{d}}}\ket{j-1}\bra{j-1}.$ | . (192)
In other words, we introduce localized dissipation to the left side of the particle [see Fig. 22(a) for a schematic illustration]. Here, we exploit the continuous quantum Zeno effect [111] in the strong-dissipation regime $\gamma_{\mathrm{d}}\gg J$ to suppress the hopping of the particle to the site that is subject to strong dissipation. Because of this feedback control of dissipation, hopping of a particle to the right is facilitated, and chiral transport is eventually achieved.
The eigenspectrum of a quantum channel for this dissipative feedback control is shown in Fig. 22(b). The eigenspectrum under the PBC forms a loop, and that under the OBC displays the non-Hermitian skin effect. Here, the loop structure of the eigenspectrum under the PBC is realized due to the dissipation since the Hamiltonian (191) does not contain a wall potential. Thus, the topological feedback control and the associated chiral transport are achieved by dissipation.
@@ X Conclusion and outlook
In this paper, we have presented a class of dynamical topological phases realized with feedback control, which are characterized by the topology of quantum channels. The nontrivial point-gap topology of a quantum channel characterized by a winding number ensures robust chiral transport, which is analogous to the chiral edge current in the quantum Hall effect and leads to the emergence of the non-Hermitian skin effect under the open boundary condition. We have also provided a symmetry classification of quantum feedback control, identified ten classes compatible with projective measurements, and constructed examples of feedback control in each symmetry class. This general framework of symmetry and topology of quantum feedback control provides a guiding principle for designing stable feedback control of quantum systems since topology is insensitive to small disturbances. Furthermore, feedback control offers a unique way to control dynamical topological phases of matter distinct from other well-known classes such as periodically driven systems [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], non-Hermitian systems [32, 33, 34, 35, 36, 37, 38], and open quantum systems subject to engineered dissipation [64, 65].
@@
Topological feedback control is expected to be realized in various experimental platforms. For example, the past decade has witnessed remarkable progress in site-resolved measurement techniques and control of ultracold atoms in optical lattices [112, 113, 114]. More recently, atoms in optical tweezer arrays have emerged as highly controllable quantum simulators in which individual atoms can be measured and addressed in a programmable manner [115, 116, 117, 118, 119, 50, 120]. The high-precision position measurement and subsequent local control of a Hamiltonian for topological feedback control can be achieved by these state-of-the-art techniques. In addition to these promising platforms, quantum feedback control has widely been used in photonic systems [42, 58], levitated nanoparticles [44, 45, 46], and superconducting qubits [43, 60, 61, 62]. Given the rapid development in quantum control, topological feedback control in various experimental platforms will expand the scope of topological phases of matter.
@@
We conclude this paper by discussing future perspectives and open problems. First, while we have mostly focused on feedback control with projective measurements, the present framework for topology of quantum channels is also applicable to general quantum measurements; therefore, it is natural to ask whether nonprojective measurements can be harnessed to construct topological feedback control. In particular, since the symmetry classification in Sec. V.3 is based on the spectral property of projective measurement channels, BL symmetry classes other than those in Table 2 are not prohibited for general quantum measurements. Since a quantum channel in a symmetry class beyond the tenfold way necessarily has a pair of eigenvalues ±$\pm 1$ (if one assumes the existence of a steady state), its eigenvalue spectrum is reminiscent of that of dissipative time crystals [33, 121], where the eigenvalue $-1$ implies oscillating components under repeated application of a quantum channel.
@@
Second, as discussed in Sec. VIII, a crucial distinction between quantum channels and general non-Hermitian matrices is complete positivity and trace preservation, which should be satisfied by any time evolution of the density matrix of a quantum system [47]. Although the topological classification of non-Hermitian matrices without the CPTP condition has been obtained [34, 35], it is unclear whether the CPTP condition alters the topological classification. For example, in Sec. VI.2, we have shown that the CPTP condition places a nontrivial constraint on the Bloch matrix with transposition symmetry, which turns out to be characterized by an integer topological invariant rather than a $\mathbb{Z}_{2}$ invariant. It merits further study to investigate the generality of this phenomenon. From a broader perspective, this problem may be related to a classification of time evolution of open quantum systems described by CPTP maps.
@@
Last but not least, topological feedback control suggests a novel connection between topology and Maxwell’s demon. For example, an experimental realization of Maxwell’s demon in Ref. [54] demonstrated chiral transport of a particle due to feedback control, in close analogy with the chiral Maxwell demon in Sec. IV (see also Ref. [82]). Maxwell’s demon has attracted century-long interest in thermodynamics and statistical physics, and an information-theoretic understanding of the role of Maxwell’s demon has led to the integration of thermodynamics and information theory [53]. Thus, investigating the topology of feedback control may shed light on a hitherto unnoticed role of topology in thermodynamics.